I have to comment on the profitability of Barclays bank. I was wondering what the best measures of profitability were (any formulas)? and where in the financial statements i could find figures for these? e.g. balance sheet/ income statement.

Many Thanks.
Answer #1
Formulae used:
Ratios to assess bank's profitability :
Return on assets(ROA)= Net income/Av. Total assets
Av. Total assets= (Beginning total assets+Ending total assets)/2
Return on equity (ROE)= Net income/Av. Total equity
Av. Total equity= (Beginning total equity+Ending total equity)/2
Net Interest Margin=Net interest income/Av. Total Assets
Website address: https://last10k.com/sec-filings/jo
(Fig. in mlns.) Fiscal yr. ended Dec 31,
2018 2017 2016
Pg. 105 1.Profit after tax 835 -1154 3729
Pg.107 2. Total assets 877700 1129343
3.Average total assets
(1129343+877700)/2= 1003521.5
4. ROA= (1/3) 835/1003522=
0.08%
Pg.107 5.Total equity 47711 65734
6.Average total equity
(65734+47711)/2= 56723
7.ROE=(1/6) 835/56723=
1.47%
Pg. 105 8.Net interest Income 3130 3876 4776
9.Net Interest margin=8/3) 3130/1003522=
0.31%
## Assets & equity are averaged
(Fig. in mlns.) Fiscal yr. ended Dec 31,
2018 2017 2016
1.Profit after tax 835 -1154 3729
2. Total assets 877700 1129343
3. ROA= (1/2) 835/877700=
0.10%
5.Total equity 47711 65734
6.ROE=(1/5) 835/47711=
1.75%
7.Net interest Income 3130 3876 4776
8.Net Interest margin=(7/2) 3130/877700=
0.36%